# Harmonics of Learning: Universal Fourier Features Emerge in Invariant Networks

### 13 Dec 2023 v1

@@ Abstract
Figure 1: Emergent Circular Harmonics. Weights learned by a neural network trained for invariance to planar rotation resemble circular haromonics. Data from [44].
@@ 1 Introduction and Related Work
Spatially localized versions of canonical 2D Fourier basis functions, such as Gabor filters or wavelets, are perhaps the most frequently observed universal features in image models. They commonly arise in the early layers of vision models – trained with efficient coding [36, 4], classification [35], temporal coherence [23], and next-step prediction [11] objectives – as well as in the primary visual cortices of diverse mammals – including cats [21], monkeys [22], and mice [10]. Non-localized Fourier features have been observed in networks trained to solve tasks that permit cyclic wraparound – for example, modular arithmetic [34], more general group compositions [7], or invariance to the group of cyclic translations [44]. In the domain of spatial navigation, the so-called grid cells of the entorhinal cortex [33] display periodic firing patterns at different spatial frequencies as they build a map of space. Their response properties are naturally modeled with the harmonics of the twisted torus [15, 37, 13]. Similar features also emerge in artificial neural networks trained to solve spatial navigation tasks [2, 8]. The ubiquity of these features across diverse learning systems is both striking and unexplained.
@@ 3 Theoretical Results
Let $G$ be a finite group, $\mathcal{H}$ be a set, and $V_{1},\cdots,V_{k}$ be complex finite-dimensional Hilbert spaces. In what follows, we will consider the space,
@@
Note that the unitarity assumption above implies the surjectivity assumption from Proposition 3.1. Therefore, it follows that for every $i$, there exists an irreducible unitary representation $\rho_{i}:\ G\rightarrow\textnormal{U}(V_{i})$ such that $W_{i}(g)=W_{i}(1)\ \rho_{i}(g)^{\dagger}$. We wish to show that if $i\not=j$ then $\rho_{i}$ and $\rho_{j}$ are non-isomorphic representations. If not, the orthogonality assumption implies that
@@
We now wish to discuss the other crucial assumption of Proposition 3.1 requiring that $\varphi(g\cdot W)=\varphi(W)$ for all $g\in G$, which is reminiscent of invariance. However, when $\mathcal{H}$ is a space of functions, we are typically interested in models that are invariant in the input variable rather than the weight variable. Therefore, we introduce the following condition, aimed at reconciling inputs and weights. To this end, suppose that $\mathcal{H}$ is a set of functions $\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{X}$ is a set acted upon by $G$ and $\mathcal{Y}$ is a set. We adhere to the notation $\varphi(W,x)=\varphi(W)(x)$, $x\in\langle G\rangle$, $W\in\mathcal{W}$, for simplicity.
@@
The adjunction property implies that if $\varphi(W,x)$ is invariant in $x$, then $\varphi(g\cdot W,x)=\varphi(W,x)$ for all $x,g$, recovering the assumption of Proposition 3.1.
@@ 3.1.2 McCulloch-Pitts Neurons and Deep Networks
Differently from Section 3, for the next result we will restrict $\mathcal{W}$ to the subspace of $\langle G\rangle\otimes\mathbb{C}^{k}$ consisting of $W$ such that the components $W_{i}$ are orthonormal. This implies, in particular, the constrain $k\leq|G|$. Note that $\mathcal{W}$ is closed by the actions of $G$ and $\textnormal{U}(\mathbb{C})$. The orthonormality condition is anyway necessary in order to recover the full Fourier transform, as stated in Corollary 3.2.
@@ 3.2 Group Recovery
Proposition 3.2 allows us to recover the group structure of $G$ up to isomorphism from the weights of a map with unitary symmetries. In other words, this enables the recovery of an unknown group in a data-driven manner from the weights of an invariant machine learning model, addressing the problem of symmetry discovery discussed in Section 1.1. The procedure was originally suggested and validated empirically in [44].
To this end, assume that $\varphi_{i}$ satisfies the requirements of Proposition 3.2. Moreover, we introduce the additional assumption that $W_{i}(1)=I\in\textnormal{End}(V_{i})$ for all $i$. If that is the case, $W$ coincides exactly with the Fourier transform by Proposition 3.2. This implies that the multiplication table of $G$ can be recovered by:
@@ Appendix B Spectral Invariants
The spectra of order $n$ are invariant polynomials of degree $n+1$ containing one conjugate variable. The presence of the latter is necessary for invariance. For $n=1,2$ they are alternatively referred to as power spectra and bispectra respectively. Note that the power spectra reduce simply to $\beta_{\rho}(x)=|\hat{x}_{\rho}|^{2}$, $\rho\in G^{\vee}$, and constitute a standard tool in signal processing. Bispectra, together with higher-order spectra, were first introduced in [27]. It is immediate to see that spectra generate all the polynomial invariants of $\langle G\rangle$ (see also [47, Theorem 2.1.4]).

### 23 Dec 2023 v2

@@ Abstract
Figure 1: Emergent Circular Harmonics. Weights learned by a neural network trained for invariance to planar rotations resemble circular harmonics. Data from [44].
@@ 1 Introduction and Related Work
Spatially localized versions of canonical 2D Fourier basis functions, such as Gabor filters or wavelets, are perhaps the most frequently observed universal features in image models. They commonly arise in the early layers of vision models – trained with efficient coding [36, 4], classification [35], temporal coherence [23], and next-step prediction [11] objectives – as well as in the primary visual cortices of diverse mammals – including cats [21], monkeys [22], and mice [10]. Non-localized Fourier features have been observed in networks trained to solve tasks that permit cyclic wraparound – for example, modular arithmetic [34], more general group compositions [7], or invariance to the group of cyclic translations [44]. In the domain of spatial navigation, the so-called grid cells of the entorhinal cortex [33] display periodic firing patterns at different spatial frequencies as they build a map of space. Their response properties are naturally modeled with the harmonics of the twisted torus [15, 37, 13]. Similar features also emerge in artificial neural networks trained to solve spatial navigation tasks [2, 8, 47]. The ubiquity of these features across diverse learning systems is both striking and unexplained.
@@ 3 Theoretical Results
Let $G$ be a finite group, $\mathcal{H}$ be a set, and $V_{1},\cdots,V_{k}$ be complex finite-dimensional Hilbert spaces. In what follows, we will consider the space
@@
Note that the unitarity assumption above implies the surjectivity assumption from Theorem 3.1. Therefore, it follows that for every $i$, there exists an irreducible unitary representation $\rho_{i}:\ G\rightarrow\textnormal{U}(V_{i})$ such that $W_{i}(g)=W_{i}(1)\ \rho_{i}(g)^{\dagger}$. We wish to show that if $i\not=j$ then $\rho_{i}$ and $\rho_{j}$ are non-isomorphic representations. If not, the orthogonality assumption implies that
@@
We now wish to discuss the other crucial assumption of Theorem 3.1 requiring that $\varphi(g\cdot W)=\varphi(W)$ for all $g\in G$, which is reminiscent of invariance. However, when $\mathcal{H}$ is a space of functions, we are typically interested in models that are invariant in the input variable rather than the weight variable. Therefore, we introduce the following condition, aimed at reconciling inputs and weights. To this end, suppose that $\mathcal{H}$ is a set of functions $\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{X}$ is a set acted upon by $G$ and $\mathcal{Y}$ is a set. We adhere to the notation $\varphi(W,x)=\varphi(W)(x)$, $x\in\langle G\rangle$, $W\in\mathcal{W}$, for simplicity.
@@
The adjunction property implies that if $\varphi(W,x)$ is invariant in $x$, then $\varphi(g\cdot W,x)=\varphi(W,x)$ for all $x,g$, recovering the assumption of Theorem 3.1.
@@ 3.1.2 McCulloch-Pitts Neurons and Deep Networks
Differently from Section 3, for the next result we will restrict $\mathcal{W}$ to the subspace of $\langle G\rangle\otimes\mathbb{C}^{k}$ consisting of $W$ such that the components $W_{i}$ are orthonormal. This implies, in particular, the constraint $k\leq|G|$. Note that $\mathcal{W}$ is closed by the actions of $G$ and $\textnormal{U}(\mathbb{C})$. The orthonormality condition is anyway necessary in order to recover the full Fourier transform, as stated in Corollary 3.2.
@@ 3.2 Group Recovery
Corollary 3.2 allows us to recover the group structure of $G$ up to isomorphism from the weights of a map with unitary symmetries. In other words, this enables the recovery of an unknown group in a data-driven manner from the weights of an invariant machine learning model, addressing the problem of symmetry discovery discussed in Section 1.1. The procedure was originally suggested and validated empirically in [44].
To this end, assume that $\varphi_{i}$ satisfies the requirements of Corollary 3.2. Moreover, we introduce the additional assumption that $W_{i}(1)=I\in\textnormal{End}(V_{i})$ for all $i$. If that is the case, $W$ coincides exactly with the Fourier transform by Corollary 3.2. This implies that the multiplication table of $G$ can be recovered by:
@@ Appendix B Spectral Invariants
The spectra of order $n$ are invariant polynomials of degree $n+1$ containing one conjugate variable. The presence of the latter is necessary for invariance. For $n=1,2$ they are alternatively referred to as power spectra and bispectra respectively. Note that the power spectra reduce simply to $\beta_{\rho}(x)=|\hat{x}_{\rho}|^{2}$, $\rho\in G^{\vee}$, and constitute a standard tool in signal processing. Bispectra, together with higher-order spectra, were first introduced in [27]. It is immediate to see that spectra generate all the polynomial invariants of $\langle G\rangle$ (see also [48, Theorem 2.1.4]).